# Multimodal Conversation Structure Understanding

### v1

@@ 1 Introduction
Multi-party conversation understanding involves identifying and structuring interactions between multiple speakers and recipients in a face-to-face setting. Such conversations typically involve complex floor-claiming and turn-taking patterns marked by various pragmatic cues. Modeling these structures is useful for any dialogue system, and it has the potential to shed light on the power dynamics between interlocutors, reveal implicit social hierarchies, and enhance our understanding of how conversational roles shift over time, among other things. While multimodal large language models (LLMs) have shown potential in reasoning over complex videos and dialogues, whether they can adequately parse the structure of conversation—resolve the reply-to relationship between utterances, or attribute roles like speakers and addressees—remains underexplored.
In this work, we draw on literature in sociolinguistics [1, 2, 3] and conversational organization [4, 5] to devise a suite of tasks for multi-modal multi-party conversation structure understanding that seeks to capture the interactional patterns of interlocuters. Building on this framework, we formally introduce a suite of tasks focused on conversational role attribution (speaker, addressees, side-participants) and conversation disentanglement (utterance linking and clustering). To support those tasks, we build on TVQA and present a human annotated dataset of 4,378 annotations for speakers and reply-to relationship between utterances, 5,599 addressees, and 3,412 side-participants. We evaluate popular vision–language models and audio-visual LLMs on this dataset, and find that there remains room for improvement for modern multimodal LLMs: while Gemini 2.0 Flash outperforms vision–language models we study, its performance degrades when conversation participants are anonymized. The superior performance of Gemini correlates with visual and acoustic clarity, which points to identity-awareness and cross-modal cues as promising directions for future work in multimodal dialogue understanding.
@@ 5.3 LLM zero-shot performance
The standard setup in video understanding often involves sampling visual frames from a video while discarding its audio. Here, we use a modified context $C_{i}=\mathcal{U}^{\star}_{i}$, and $\mathcal{U}_{i}=\{(v^{\star}_{j},u_{j},\;f_{j})\}_{j=1}^{|\mathcal{U}|}$: $v^{\star}_{i}$ is the final frame sampled from $V_{i}$ in the interval $[t_{s,j}\,,t_{e,j}]$, interleaving with the words of $u_{j}$; this is further augmented with audio-derived speaker features $f$ extracted using Librosa [51]: speech rate (slower or faster), pitch range (wide or narrow), pitch mean (lower or higher), and spectral centroid (darker or brighter), which collectively aim to capture the acoustic profile of the speaker.
@@ A.1.4 Annotation quality
We report inter-annotator agreement in Table 7, using the same set of evaluation metrics applied to our models. Agreement is computed as the average of all pairwise comparisons among the four primary annotators. We observe a clear improvement across all metrics in the second round of annotation, which reflects increased consistency and agreement in annotation after iterative refinement.

### v2

@@ 1 Introduction
Multi-party conversation understanding involves identifying and structuring interactions between multiple speakers and recipients in a face-to-face setting. Such conversations typically involve complex floor-claiming and turn-taking patterns marked by various pragmatic cues. Modeling these structures is useful for any dialogue system, and it has the potential to shed light on the power dynamics between interlocutors, reveal implicit social hierarchies, and enhance our understanding of how conversational roles shift over time, among other things. While multimodal large language models (LLMs) have shown potential in reasoning over complex videos and dialogues, whether they can adequately parse the structure of conversation—resolve the reply-to relationship between utterances, or attribute roles like speakers and addressees—remains underexplored. To bridge this gap, in this work we make the following contributions:
Figure 1: Our proposed structured prediction task for multimodal conversation structure understanding. Grounded in sociolinguistic and conversational analysis [1, 2, 3, 4, 5], the task requires predicting, for each utterance in the given clip: the speaker, addressee(s), side-participants, and the utterance it replies to. The example, taken from the Big Bang Theory in TVQA [28, 29], illustrates our unified formulation, which treats conversational role attribution and conversation disentanglement as complementary subtasks for modeling the interactional dynamics of dialogue. Further analysis of this example can be found in §A.1.3.
  • We draw on literature in sociolinguistics [1, 2, 3] and conversational organization [4, 5] to devise a framework for multimodal multi-party conversation structure understanding that seeks to capture the interactional patterns of interlocuters. Building on this framework, we formally introduce a suite of tasks (Fig. 1) focused on conversational role attribution (speaker, addressees, side-participants) and conversation disentanglement (utterance linking and clustering).
  • To support those tasks, we build on TVQA [28, 29] and present TV-MMPC, a human-annotated dataset of 4,378 annotations for speakers and reply-to relationship between utterances, 5,599 addressees, and 3,412 side-participants, available at https://doi.org/10.7910/DVN/4KUKUL.
  • We evaluate popular vision–language models and audio-visual LLMs on this annotated dataset, and find that there remains room for improvement for modern multimodal LLMs: while Gemini 2.0 Flash outperforms vision–language models we study, its performance degrades when conversation participants are anonymized; its superior performance correlates with visual and acoustic clarity. This points to memorization and cross-modal cues, among others, as promising directions for future work in multimodal dialogue understanding.
@@ 5.3 LLM zero-shot performance
The standard setup in video understanding often involves sampling visual frames from a video while discarding its audio. Here, context $C_{i}=\mathcal{U}^{\star}_{i}$, and $\mathcal{U}_{i}=\{(v^{\star}_{j},u_{j},\;f_{j})\}_{j=1}^{|\mathcal{U}|}$: $v^{\star}_{i}$ is the final frame sampled from $V_{i}$ in the interval $[t_{s,j}\,,t_{e,j}]$, interleaving with the words of $u_{j}$; this is further augmented with audio-derived speaker features $f$ extracted using Librosa [51]: speech rate (slower or faster), pitch range (wide or narrow), pitch mean (lower or higher), and spectral centroid (darker or brighter), which collectively aim to capture the acoustic profile of the speaker.
@@ A.1.4 Annotation quality
We report inter-annotator agreement in Table 7 based on 10–12 sampled clips (not part of the final dataset). Agreement is computed as the average of all pairwise comparisons among the four primary annotators, using the same set of evaluation metrics for experiments, and we observe a clear improvement in the main round of annotation.